# Experimental demonstration of spontaneous symmetry breaking with emergent multi-qubit entanglement

### v2

@@ Abstract
The most typical example for illustrating SSB is a classical model, where a point-like particle moves along a straight line under the influence of a sombrero potential, given by $V(x)=\Lambda x^{4}-\mu x^{2}$ with $\Lambda,\mu>0$ [12]. The particle has two different positions of stable equilibrium, $x=\pm\sqrt{\mu/2\Lambda}$. Although the potential possesses a global spatial-inversion symmetry, i.e., $V(-x)=V(x)$, each of the two equilibrium positions loses this symmetry. Which equilibrium position the particle chooses to stay at depends on perturbation. This is the standard understanding of SSB, which, however, does not grasp the quantum-mechanical feature of SSB. For quantum systems, the superposition principle allows degenerate stable symmetry-breaking eigenstates to be superimposed with each other, forming a symmetric one. As such, SSB is said to occur if there exist two or more degenerate eigenstates, each of which lacks the symmetry of the governing Hamiltonian [12]. When the system initially has the same symmetry as the Hamiltonian, it should evolve to the symmetric superposition of the corresponding degenerate eigenstates after a quantum phase transition under the Hamiltonian dynamics. For an open quantum system, it is the environment-induced decoherence that destroys the quantum coherence, forcing the system to choose one of the symmetry-breaking eigenstates as the reality [13, 14, 15].
@@
During the past few years, quantum phase transitions have been probed in different many-body systems, including neutral atom arrays [16, 17], trapped ions [18, 19, 20, 21], superconducting qubits [22, 23, 24], and ultracold atoms trapped in optical cavities [25, 26, 27, 28, 29, 30, 31]. In particular, symmetry-breaking phases have been experimentally explored in Dicke-like quantum phase transitions [26, 27, 28]. However, the driven-dissipative character of the realized phase transitions makes it impossible to observe the quantum coherence between the SSB degenerate eigenstates. So far, experimental demonstrations of SSB with quantum-mechanical features have been confined to a photonic mode that is coupled to a single qubit [32, 33]. For a finite many-qubit system, one of the most essential characteristics of the SSB phase transition is the appearance of a multi-qubit maximally entangled state [34, 35, 36], also referred to as the Greenberger-Horne-Zeilinger (GHZ) state [37]. Despite being fundamentally important, such a highly nonclassical SSB feature has not been experimentally observed.
@@
We here demonstrate the purely quantum-mechanical SSB, featuring emergent GHZ entanglement, in the multi-qubit Lipkin-Meshkov-Glick (LMG) model [38], governed by a Hamiltonian featuring the competition between individual continuous drivings and intra-qubit interactions. The LMG model is synthesized in a superconducting circuit, where 6 Xmon qubits are nearly homogeneously coupled to one another with the assistance of a bus resonator connected to all these qubits. By slowly decreasing the driving strength, we realize the SSB process, during which the multi-qubit system evolves from an initial symmetric product state to a superposition of two degenerate symmetry-breaking eigenstates of the governing Hamiltonian. Through the real-time probing of the average two-qubit longitudinal correlation during the quench process, we observe the progressive improvement of the $Z_{2}$ symmetry breaking degree. We verify the quantum character of the SSB by measuring multi-qubit transverse quantum correlations and mapping out the Wigner function. Our results reveal the fundamental distinction between the SSB of quantum many-body systems and that of classical ones, and offer a dephasing-insensitive strategy for multi-qubit entanglement engineering in solid-state systems.
@@
where $\Omega$ is the driving strength, $\lambda$ denotes the qubit-qubit coupling, $S_{x/z}=\frac{1}{2}\stackrel{{\scriptstyle N}}{{\mathrel{\mathop{\sum}\limits_{j=1}}}}\sigma_{j}^{x/z}$ is the collective angular operator, and $\sigma_{j}^{x/z}$ denotes the Pauli operators of the $j$th qubit. The Hamiltonian has the $Z_{2}$ symmetry, defined as the joint parity: ${\cal P}_{x}=\otimes_{j=1}^{N}\sigma_{j}^{x}$. The system’s ground-state behavior depends on the competition of longitudinal couplings and transverse drivings. When $\Omega,\lambda>0$ and $\Omega\gg\lambda$, the longitudinal couplings are suppressed, so that the system presents a unique ground state of the form $\left|X\right\rangle=\otimes_{j=1}^{N}\left|+_{j}\right\rangle$, where |±$\left|\pm_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle\pm\left|e_{j}\right\rangle)$. Here $\left|g_{j}\right\rangle$ and $\left|e_{j}\right\rangle$ denote the lower and upper levels of the $j$th qubit. This eigenstate has the same symmetry as the Hamiltonian, i.e., ${\cal P}_{x}\left|X\right\rangle=\left|X\right\rangle$. When $\Omega\ll\lambda$, the transverse driving only slightly shifts the energy levels of the eigenstates $\left|J,m\right\rangle$ ($\left|J,m\right\rangle$ is the eigenstate of $S_{z}$ with eigenvalue $m$ and total angular momentum quantum number $J$. See Supplementary Material [39] Sec. S1 for the detailed derivation). The most remarkable feature of the system in the strong-coupling regime is that it has two degenerate ground states, $\left|J,J\right\rangle=\otimes_{j=1}^{N}\left|e_{j}\right\rangle$ and $\left|J,-J\right\rangle=\otimes_{j=1}^{N}\left|g_{j}\right\rangle$, each breaking the system’s $Z_{2}$ symmetry.
@@
Suppose that the system is initially in the state $\left|X\right\rangle$, which is the lowest eigenstate of the Hamiltonian under the condition $\lambda/\Omega\rightarrow 0$. When the control parameter $\Omega$ is slowly decreased, the SSB will occur. In the limit $\Omega\ll\lambda$, the system will be moved to the symmetric superposition of the two degenerate ground states, $\left|\psi_{+}\right\rangle=\frac{1}{\sqrt{2}}(\left|J,J\right\rangle+\left|J,-J\right\rangle)$, which corresponds to an $N$-qubit GHZ state. As the slow variation of the transverse drive does not change the symmetry of the Hamiltonian, the simultaneous emergence of the two symmetry-breaking eigenstates is a result of SSB quantum phase transition. The SSB is manifested by the vulnerable entanglement properties of the GHZ state, which will be transformed into a classical mixture by the environment-induced decoherence within a timescale $\tau_{\rm dec}\sim 1/N$ [15]. We note that the adiabatic preparation of superpositions of two degenerate ground states has been demonstrated in Rydberg atom arrays [40], but where the emergence of the degeneracy cannot be attributed to SSB as the Hamiltonian itself has no symmetry.
@@
Figure 1: Experimental realization of SSB in a multi-qubit system. (a) Device schematic. The device comprises 10 Josephson-junction-based qubits and a bus resonator. The first 6 qubits, denoted as $Q_{j}$ ($j=1$ to $6$), are approximately homogeneously coupled to the resonator with the on-resonance photonic swapping rate $\xi\simeq 2\pi\times 20$ MHz. The other four qubits are unused in the experiment (not shown here). (b) Synthesis of the LMG model. When these qubits are detuned from the resonator by the same amount $|\Delta|\gg\xi$, any pair of qubits, $Q_{j}$ and $Q_{k}$, are coupled through virtual photon exchange mediated by the resonator, with the effective coupling strength $\lambda\simeq\xi^{2}/|\Delta|$. Each of these qubits is driven by a continuous microwave with the Rabi frequency $\Omega$. The system ground state depends upon the competition between resonator-induced qubit-qubit couplings and transverse drivings. (c) Pulse sequence. The experiment starts by tuning the qubits to the corresponding idle frequencies, where they are prepared in the state $\otimes_{j=1}^{6}\left|+_{j}\right\rangle$ with a continuous microwave pulse. By tuning all the qubits to the same operation frequency and applying continuous drivings, the system dynamics is described by the LMG Hamiltonian. The subsequent quench process is realized by modeling the driving amplitudes as $\Omega(\tau)=\Omega^{0}\exp(-\tau/t_{f})$, where $\Omega^{0}/(2\pi)=40$ MHz, $t_{f}=60$ ns, $\tau$ represents the quench time. Following a preset quench process, the qubits are biased back to their idle frequencies for the state readout.
@@
Our circuit QED system is composed of 10 frequency-tunable superconducting qubits ($Q_{j}$), coupled to a bus resonator with a fixed frequency $\omega_{b}/(2\pi)\simeq 5.796$ GHz. The LMG model is realized with the first 6 qubits that are almost homogeneously coupled to the resonator with the photonic swapping rate $\xi\simeq 2\pi\times 20$ MHz, as sketched in Fig. 1(a). The other 4 unused qubits (not shown) are highly detuned from the resonator and the first 6 qubits, so that they do not affect the engineered LMG model. The energy decaying times ($T_{1}$) of these qubits range from 14 to 28 $\mu$s, while the dephasing times ($T_{2}$), measured at the corresponding idle frequencies, range from 2 to 5 $\mu$s. The system parameters are detailed in the Supplementary Material [39] Sec. S2 (see also Ref. [44]). As the qubits’ maximum frequencies in our system are close to $\omega_{b}$, the LMG model is synthesized at the operation frequency $\omega_{o}/(2\pi)=5.6895$ GHz, which is lower than $\omega_{b}/(2\pi)$ by an amount of 106.5 MHz [Fig. 1(b)]. The resulting intra-qubit couplings induced by the virtual photon exchange have a magnitude of $\lambda=2\pi\times 3.8$ MHz. The combination of these effective swapping couplings and transverse drivings with a zero phase is described by the effective Hamiltonian $-H$. The SSB quantum phase transition associated with the highest eigenstate of $-H$ is equivalent to the ground phase transition of $H$. Before the experiment, all qubits are initialized to their ground states at the idle frequencies. The pulse sequence for realizing the quantum phase transition is shown in Fig. 1(c). The experiment starts by tuning all these qubits to their idle frequencies. Then a $\pi/2$ pulse is applied to each qubit, transforming it to the superposition state $\left|+_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle+\left|e_{j}\right\rangle)$. This is followed by tuning all the qubits to the operation frequency and applying continuous drivings, incorporating the transverse drivings with the longitudinal couplings. The phase transition is realized by a quench process, during which the driving strengths are modeled as $\Omega(\tau)=\Omega^{0}\exp(-\tau/t_{f})$, where $\Omega^{0}/(2\pi)=40$ MHz, $t_{f}=60$ ns, and $\tau$ represents the quench time, as shown in Fig. 2(a). After a preset quench time, the transverse drives are switched off, following which each of the qubits is biased back to its idle frequency for the state readout.
@@
Figure 2: Observed populations evolutions. (a) $\Omega$ versus $\tau$ during the quench process. (b) Measured population, $P_{0}^{e}$, as a function of the quench time. (c), (d), (e), (f) Population distributions measured at $\tau=0$ ns (c), $\tau=50$ ns (d), $\tau=100$ ns (e), and $\tau=150$ ns (f). For clarity, we combine all populations with the same excitation number $n$ into a single one, denoted as $P_{n}^{e}$.
@@
To characterize the state evolution, it is necessary to be able to distinguish between $\left|\psi_{0}\right\rangle=\left|+_{j}\right\rangle^{\otimes 6}$ and $\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle^{\otimes 6}+\left|e_{j}\right\rangle^{\otimes 6})$ that are the ground states in the limits $\lambda/\Omega\rightarrow 0$ and $\Omega/\lambda\rightarrow 0$, respectively. Though both states possess a zero z-axis magnetization $\sum_{j=1}^{N}\left\langle\sigma_{j}^{z}\right\rangle$, they have very different population distributions. For $\left|\psi_{0}\right\rangle$, the probability of all qubits being populated in $\left|g\right\rangle$ is $P_{0}^{e}=1/2^{6}$, while that for $\left|\psi_{1}\right\rangle$ is $1/2$. The measured population, $P_{0}^{e}$, as a function of the quench time, is displayed in Fig. 2(b). As expected, $P_{0}^{e}$ is increased monotonously with the quench process, during which $\Omega/\lambda$ is decreased. To reveal the system evolution more detailedly, we display the population distributions measured at $\tau=0$, $\tau=50$, $\tau=100$, and $\tau=150$ ns in Figs. 2(c), 2(d), 2(e) and 2(f), respetively, which confirm the system population undergoes a significant change during the quench, transformed from a binomial distribution to a pattern dominated by the two maximally distinct components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. Each of these components corresponds to a symmetry-breaking order with respect to the z-axis magnetization. We note that the dynamics of the system deviates from a unitary process due to dissipation.
@@
${\cal C}_{6}^{TQ}=\left\langle(\sigma_{j}^{\beta})^{\otimes 6}\right\rangle-\left\langle\sigma_{j}^{\beta}\right\rangle^{\otimes 6},$ (2)
@@
where $\sigma_{j}^{\beta}=\cos\beta\sigma_{j}^{x}+\sin\beta\sigma_{j}^{y}$. To measure ${\cal C}_{6}^{TQ}$, before the state readout, each qubit is subjected to a $\pi/2$ rotation around the axis with an angle $\beta-\pi/2$ to the x-direction on the equatorial plane. The off-diagonal element of the system density matrix, $\rho_{e,...,e;g,...,g}$, is directly related to ${\cal C}_{6}^{TQ}$, by ${\cal C}_{6}^{TQ}=2\left|\rho_{e,...,e;g,...,g}\right|\cos(6\beta-\gamma)$, where $\gamma=\arg\rho_{e,...,e;g,...,g}$ [44, 46]. Figure 3(a) shows ${\cal C}_{6}^{TQ}$ versus $\beta$, measured for different quench times, from which the visibility of the measured ${\cal C}_{6}^{TQ}$ as a function of $\tau$ is inferred and shown in Fig. 3(b). The results show that the quantum coherence $\left|\rho_{e,...,e;g,...,g}\right|$ does not show a monotonous behavior during the quench, which can be interpreted as a consequence of the competition between the coherent evolution and the relaxation. The Hamiltonian dynamics renders the system population to converge to $\left|g_{j}\right\rangle^{\otimes N}$ and $\left|e_{j}\right\rangle^{\otimes N}$, while the relaxation deteriorates their quantum coherence. At the beginning, the populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$ are almost 00, and so is their coherence. As the amount of coherence lost during a given time is proportional to the coherence itself, at the early stage the Hamiltonian dynamics dominates the relaxation. With the increase of the populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, the relaxation plays a more and more important role. When the decoherence-induced coherence loss exceeds the gain due to the Hamiltonian dynamics, ${\cal C}_{6}^{TQ}$ begins to descend. The sources of experimental errors include decoherence, nonadiabaticity, imperfect preparation of the initial symmetric state $\left|X\right\rangle$, limitation of the qubits-resonator detuning, and the nonhomogeneous couplings (see Supplemental Material [39] Sec. S4 for details). As dephasing noises are suppressed by the transverse drivings [47], the energy relaxation is the main decoherence effect.
@@
Figure 3: Measured quantum correlations. (a) 6-qubit transverse quantum correlation ${\cal C}_{6}^{TQ}$ versus $\beta$, measured for quench times $\tau=0$, 75, 105, and 150 ns. Here ${\cal C}_{6}^{TQ}$ is defined as $\left\langle\otimes_{j=1}^{6}\sigma_{j}^{\beta}\right\rangle-\left\langle\sigma_{j}^{\beta}\right\rangle^{\otimes 6}$, where $\sigma_{j}^{\beta}$ is the spin component along the axis with an angle $\beta$ to the x-axis on the equatorial plane of the Bloch sphere. (b) Fringe visibility of the measured ${\cal C}_{6}^{TQ}$ as a function of $\tau$. (c) Average 2-qubit longitudinal correlation ${\cal C}_{2}^{L}$ versus $\tau$.
@@
This transverse quantum correlation can measure the extent to which the symmetry is broken only when there are no decoherence effects, as it is zero for any classical mixture of the two symmetry-breaking components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. To quantify the symmetry-breaking extent for a general multi-qubit state, we use the average two-qubit longitudinal correlation, defined as

### v3

@@ Abstract
The most typical example for illustrating SSB is a classical model, where a point-like particle moves along a straight line under the influence of a sombrero potential, given by $V(x)=\Lambda x^{4}-\mu x^{2}$ with $\Lambda,\mu>0$ [12]. The particle has two different positions of stable equilibrium, $x=\pm\sqrt{\mu/2\Lambda}$. Although the potential possesses a global spatial-inversion symmetry, i.e., $V(-x)=V(x)$, each of the two equilibrium positions loses this symmetry. Which equilibrium position the particle chooses to stay at depends on perturbation. This is the standard understanding of SSB, which, however, does not grasp the quantum-mechanical feature of SSB. For quantum systems, the superposition principle allows degenerate stable symmetry-breaking eigenstates to be superimposed with each other, forming a symmetric one. As such, SSB is said to occur if there exist two or more degenerate eigenstates, each of which lacks the symmetry of the governing Hamiltonian [12]. When the system initially has the same symmetry as the Hamiltonian, it should evolve to the symmetric superposition of the corresponding degenerate eigenstates after a quantum phase transition (QPT) under the Hamiltonian dynamics. For an open quantum system, it is the environment-induced decoherence that destroys the quantum coherence, forcing the system to choose one of the symmetry-breaking eigenstates as the reality [13, 14, 15].
@@
During the past few years, QPT have been probed in different many-body systems, including neutral atom arrays [16, 17], trapped ions [18, 19, 20, 21], superconducting qubits [22, 23, 24], and ultracold atoms trapped in optical cavities [25, 26, 27, 28, 29, 30, 31]. In particular, symmetry-breaking phases have been experimentally explored in Dicke-like QPTS [26, 27, 28]. However, the driven-dissipative character of the realized phase transitions makes it impossible to observe the quantum coherence between the SSB degenerate eigenstates. So far, experimental demonstrations of SSB with quantum-mechanical features have been confined to a photonic mode that is coupled to a single qubit [32, 33]. For a finite many-qubit system, one of the most essential characteristics of the SSB phase transition is the appearance of a multi-qubit maximally entangled state [34, 35, 36], also referred to as the Greenberger-Horne-Zeilinger (GHZ) state [37]. Although quantum entanglement has been produced by adiabatic control of the quantum Ising Hamiltonian in a two-qubit ion trap experiment [38], such a highly nonclassical SSB feature has not been experimentally observed in QPTs of many-body systems.
@@
We here demonstrate the purely quantum-mechanical SSB, featuring emergent GHZ entanglement, in the QPT of the multi-qubit Lipkin-Meshkov-Glick (LMG) model [39, 40, 41, 42, 43], governed by a Hamiltonian featuring the competition between individual continuous drivings and intra-qubit interactions. The LMG model is synthesized in a superconducting circuit, where 6 Xmon qubits are nearly homogeneously coupled to one another with the assistance of a bus resonator connected to all these qubits. By slowly decreasing the driving strength, we realize the SSB process, during which the multi-qubit system evolves from an initial symmetric product state to a superposition of two degenerate symmetry-breaking eigenstates of the governing Hamiltonian. Through the real-time probing of the average two-qubit longitudinal correlation during the quasi-adiabatic process, we observe the progressive improvement of the $Z_{2}$ symmetry breaking degree. We verify the quantum character of the SSB by measuring multi-qubit transverse quantum correlations and mapping out the Wigner function. Our results reveal the fundamental distinction between the SSB in QPT of many-body systems and that in classical phase transitions, and offer a dephasing-insensitive strategy for multi-qubit entanglement engineering in solid-state systems, thanks to the dynamical decoupling effects produced by drivings individually applied to the qubits [44].
@@
where $\Omega$ is the driving strength, $\lambda$ denotes the qubit-qubit coupling, $S_{x/z}=\frac{1}{2}\stackrel{{\scriptstyle N}}{{\mathrel{\mathop{\sum}\limits_{j=1}}}}\sigma_{j}^{x/z}$ is the collective angular operator, and $\sigma_{j}^{x/z}$ denotes the Pauli operators of the $j$th qubit. The Hamiltonian has the $Z_{2}$ symmetry, defined as the joint parity: ${\cal P}_{x}=\otimes_{j=1}^{N}\sigma_{j}^{x}$. The system’s ground-state behavior depends on the competition of longitudinal couplings and transverse drivings. When $\Omega,\lambda>0$ and $\eta=N\lambda/\Omega\ll 1$, the longitudinal couplings are suppressed, so that the system presents a unique ground state of the form $\left|X\right\rangle=\otimes_{j=1}^{N}\left|+_{j}\right\rangle$, where |±$\left|\pm_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle\pm\left|e_{j}\right\rangle)$. Here $\left|g_{j}\right\rangle$ and $\left|e_{j}\right\rangle$ denote the lower and upper levels of the $j$th qubit. This eigenstate has the same symmetry as the Hamiltonian, i.e., ${\cal P}_{x}\left|X\right\rangle=\left|X\right\rangle$. When $\eta\gg 1$, the transverse driving only slightly shifts the energy levels of the eigenstates $\left|J,m\right\rangle$ ($\left|J,m\right\rangle$ is the eigenstate of $S_{z}$ with eigenvalue $m$ and total angular momentum quantum number $J$. See Sec. S1 of the Supplementary Material [45] for the detailed derivation). The most remarkable feature of the system in the strong-coupling regime is that it has two degenerate ground states, $\left|J,J\right\rangle=\otimes_{j=1}^{N}\left|e_{j}\right\rangle$ and $\left|J,-J\right\rangle=\otimes_{j=1}^{N}\left|g_{j}\right\rangle$, each breaking the system’s $Z_{2}$ symmetry.
@@
Suppose that the system is initially in the state $\left|X\right\rangle$, which is the lowest eigenstate of the Hamiltonian under the condition $\eta\ll 1$. When the control parameter $\Omega$ is slowly decreased, the SSB will occur. In the limit $\eta\gg 1$, the system will be moved to the symmetric superposition of the two degenerate ground states, $\left|\psi_{+}\right\rangle=\frac{1}{\sqrt{2}}(\left|J,J\right\rangle+\left|J,-J\right\rangle)$, which corresponds to an $N$-qubit GHZ state. As the slow variation of the transverse drive does not change the symmetry of the Hamiltonian, the simultaneous emergence of the two symmetry-breaking eigenstates is a result of SSB QPT. When $N$ is sufficiently large, the QPT occurs at the critical point $\eta=1$. For the six-qubit case, the system’s state is not suddenly changed at this point, but the GHZ entanglement still emerges when $\eta\gg 1$ as a consequence of SSB (see Sec. S2 of the Supplemental Material [45] for details). The SSB is manifested by the vulnerable entanglement properties of the GHZ state, which will be transformed into a classical mixture by the environment-induced decoherence within a timescale $\tau_{\rm dec}\sim 1/N$ [15]. We note that the adiabatic preparation of superpositions of two degenerate ground states has been demonstrated in Rydberg atom arrays [46], but where the emergence of the degeneracy cannot be attributed to SSB as the Hamiltonian itself has no symmetry.
@@
Figure 1: Experimental realization of SSB in a multi-qubit system. (a) Device schematic. The device comprises 10 Josephson-junction-based qubits and a bus resonator. The first 6 qubits, denoted as $Q_{j}$ ($j=1$ to $6$), are approximately homogeneously coupled to the resonator with the on-resonance photonic swapping rate $\xi\simeq 2\pi\times 20$ MHz. The other four qubits are unused in the experiment. (b) Synthesis of the LMG model. When these qubits are detuned from the resonator by the same amount $|\Delta|\gg\xi$, any pair of qubits, $Q_{j}$ and $Q_{k}$, are coupled through virtual photon exchange mediated by the resonator, with the effective coupling strength $\lambda\simeq\xi^{2}/|\Delta|$. Each of these qubits is driven by a continuous microwave with the Rabi frequency $\Omega$. The system ground state depends upon the competition between resonator-induced qubit-qubit couplings and transverse drivings. (c) Pulse sequence. The experiment starts by tuning the qubits to the corresponding idle frequencies, where they are prepared in the state $\otimes_{j=1}^{6}\left|+_{j}\right\rangle$ with a continuous microwave pulse. By tuning all the qubits to the same operation frequency and applying continuous drivings, the system dynamics is described by the LMG Hamiltonian. Following a preset quasi-adiabatic process, the qubits are biased back to their idle frequencies for the state readout.
@@
Our circuit QED system is composed of 10 frequency-tunable superconducting qubits ($Q_{j}$), coupled to a bus resonator with a fixed frequency $\omega_{b}/(2\pi)\simeq 5.796$ GHz. The LMG model is realized with the first 6 qubits that are almost homogeneously coupled to the resonator with the photonic swapping rate $\xi\simeq 2\pi\times 20$ MHz, as sketched in Fig. 1(a). The other 4 unused qubits (not shown) are highly detuned from the resonator and the first 6 qubits, so that they do not affect the engineered LMG model. The energy decaying times ($T_{1}$) of these qubits range from 9 to 32 $\mu$s, while the dephasing times ($T_{2}$) range from 2 to 5 $\mu$s. The system parameters are detailed in Sec. S3 of the Supplementary Material [45] (see also Ref. [50]). As the qubits’ maximum frequencies in our system are close to $\omega_{b}$, the LMG model is synthesized at the operation frequency $\omega_{o}/(2\pi)=5.6895$ GHz, which is lower than $\omega_{b}/(2\pi)$ by an amount of 106.5 MHz [Fig. 1(b)]. The resulting intra-qubit couplings induced by the virtual photon exchange have a magnitude of $\lambda=2\pi\times 3.8$ MHz. The combination of these effective swapping couplings and transverse drivings with a zero phase is described by the effective Hamiltonian $-H$. The SSB QPT associated with the highest eigenstate of $-H$ is equivalent to the ground phase transition of $H$. Before the experiment, all qubits are initialized to their ground states at the idle frequencies. The pulse sequence for realizing the QPT is shown in Fig. 1(c). The experiment starts by tuning all these qubits to their idle frequencies. Then a $\pi/2$ pulse is applied to each qubit, transforming it to the superposition state $\left|+_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle+\left|e_{j}\right\rangle)$. This is followed by tuning all the qubits to the operation frequency and applying continuous drivings, incorporating the transverse drivings with the longitudinal couplings. The phase transition is realized by a quasi-adiabatic process shown in Fig. 2(a). As shown in Sec. S5 of the Supplemental Material [45], the adiabatic condition can be well satisfied during this process. Then each qubit is biased back to its idle frequency for the state readout.
@@
Figure 2: Observed populations evolutions. (a) $\Omega$ versus $\tau$ during the quasi-adiabatic process, varied as $\Omega(\tau)=\Omega^{0}\exp(-\tau/t_{f})$, where $\Omega^{0}/(2\pi)=40$ MHz, $t_{f}=60$ ns, and $\tau$ represents the time. (b) Measured population, $P_{e_{1}...e_{6}}$, as a function of the time. (c), (d), (e), (f) Population distributions measured at $\tau=0$ ns (c), $\tau=50$ ns (d), $\tau=100$ ns (e), and $\tau=149$ ns (f). For clarity, we combine all populations with the same excitation number $n$ into a single one.
@@
To characterize the state evolution, it is necessary to be able to distinguish between $\left|\psi_{0}\right\rangle=\left|+_{j}\right\rangle^{\otimes 6}$ and $\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle^{\otimes 6}+\left|e_{j}\right\rangle^{\otimes 6})$ that are highest eigenstates in the limits $\eta\ll 1$ and $\eta\gg 1$, respectively. Though both states possess a zero z-axis magnetization $\sum_{j=1}^{N}\left\langle\sigma_{j}^{z}\right\rangle$, they have very different population distributions. For $\left|\psi_{0}\right\rangle$, the probability of all qubits being populated in $\left|e\right\rangle$ is $P_{e_{1}...e_{6}}=1/2^{6}$, while that for $\left|\psi_{1}\right\rangle$ is $1/2$. The measured population, $P_{e_{1}...e_{6}}$, as a function of the time, is displayed in Fig. 2(b). As expected, $P_{e_{1}...e_{6}}$ is increased monotonously with the increase of $\eta$ before $\tau=149$ ns. The progressive decrease of this population after $149$ ns is due to the energy relaxations of the qubits. To reveal the system evolution more detailedly, we display the population distributions measured at $\tau=0$, $\tau=50$, $\tau=100$, and $\tau=149$ ns in Figs. 2(c), 2(d), 2(e) and 2(f), respectively, which confirm the system population undergoes a significant change during the quasi-adiabatic process, transformed from a binomial distribution to a pattern dominated by the two maximally distinct components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. Each of these components corresponds to a symmetry-breaking order with respect to the z-axis magnetization. Due to the environment-induced decays, $\left|g_{j}\right\rangle^{\otimes 6}$ is chosen as the classical reality after a long time dynamics.
@@
${\cal C}_{6}^{T}=\left\langle(\sigma_{j}^{\beta})^{\otimes 6}\right\rangle-\left\langle\sigma_{j}^{\beta}\right\rangle^{\otimes 6},$ (2)
@@
where $\sigma_{j}^{\beta}=\cos\beta\sigma_{j}^{x}+\sin\beta\sigma_{j}^{y}$. To measure ${\cal C}_{6}^{T}$, before the state readout, each qubit is subjected to a $\pi/2$ rotation around the axis with an angle $\beta-\pi/2$ to the x-direction on the equatorial plane. The off-diagonal element of the system density matrix, $\rho_{e,...,e;g,...,g}$, is directly related to ${\cal C}_{6}^{T}$, by ${\cal C}_{6}^{T}=2\left|\rho_{e,...,e;g,...,g}\right|\cos(6\beta-\gamma)$, where $\gamma=\arg\rho_{e,...,e;g,...,g}$ [50, 52]. Figure 3(a) shows ${\cal C}_{6}^{T}$ versus $\beta$, measured for different times, from which the visibility of the measured ${\cal C}_{6}^{T}$ as a function of $\tau$ is inferred and shown in Fig. 3(b). The results show that the quantum coherence $\left|\rho_{e,...,e;g,...,g}\right|$ does not show a monotonous behavior during the quasi-adiabatic process, which can be interpreted as a consequence of the competition between the coherent evolution and the relaxation. The Hamiltonian dynamics renders the system population to converge to $\left|g_{j}\right\rangle^{\otimes N}$ and $\left|e_{j}\right\rangle^{\otimes N}$, while the relaxation deteriorates their quantum coherence. At the beginning, the populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$ are almost 00, and so is their coherence. As the amount of coherence lost during a given time is proportional to the coherence itself, at the early stage the Hamiltonian dynamics dominates the relaxation. With the increase of the populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, the relaxation plays a more and more important role. When the decoherence-induced coherence loss exceeds the gain due to the Hamiltonian dynamics, ${\cal C}_{6}^{T}$ begins to descend, and tends to 0 after a long time evolution as a consequence of decoherence. The maximum fringe contrast of ${\cal C}_{6}^{T}$ is 0.75, which, together with the measured populations of $\left|g_{j}\right\rangle^{\otimes 6}$ (0.38) and $\left|e_{j}\right\rangle^{\otimes 6}$ (0.41), yields a fidelity of 0.77 with respect to the ideal 6-qubit GHZ state. The sources of experimental errors include the nonadiabatic effect, qubits’ dephasings and energy relaxations, inhomogeneity of the intra-qubit couplings, and imperfections in qubit control. The errors caused by these imperfections are individually quantified in Sec. S6 of the Supplemental Material [45].
@@
Figure 3: Measured quantum correlations. (a) 6-qubit transverse quantum correlation ${\cal C}_{6}^{T}$ versus $\beta$, measured for times $\tau=0$, 75, 149, and 1000 ns. Here ${\cal C}_{6}^{T}$ is defined as $\left\langle\otimes_{j=1}^{6}\sigma_{j}^{\beta}\right\rangle-\left\langle\sigma_{j}^{\beta}\right\rangle^{\otimes 6}$, where $\sigma_{j}^{\beta}$ is the spin component along the axis with an angle $\beta$ to the x-axis on the equatorial plane of the Bloch sphere. (b) Fringe visibility of the measured ${\cal C}_{6}^{T}$ as a function of $\tau$. (c) Average 2-qubit longitudinal correlation ${\cal C}_{2}^{L}$ versus $\tau$.
The emergent GHZ state can exhibit multipartite and high-order quantum correlations beyond two-qubit entanglement [37, 42, 53, 43]. Furthermore, such a state only has a 1/32 overlapping with the initial state, which reflects a drastic change of the system state during the adiabatic evolution. In distinct contrast, for the two-qubit case the overlapping between the initial and final state is as high as 0.5 [38]. Recently, multi-qubit ground states of Ising models were quasi-adiabatically prepared with trapped-ion simulators [54, 55], but where multipartite quantum correlations, which represent the most remarkable nonclassical feature of SSB quantum phase transitions, were not measured.
@@
The transverse quantum correlation can measure the extent to which the symmetry is broken only when there are no decoherence effects, as it is zero for any classical mixture of the two symmetry-breaking components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. To quantify the symmetry-breaking extent for a general multi-qubit state, we use the average two-qubit longitudinal correlation, defined as